OpenStudy (anonymous):

You are to receive a 9% increase in wages per hour. If you currently make $18.55 per hour, what will your hourly wage be after the increase? Round to the nearest cent.

OpenStudy (anonymous):

Just want to make sure that I have the correct answer?

OpenStudy (tkhunny):

Please present your correct answer. An example of how you managd that answer would also be helpful.

OpenStudy (anonymous):

I multiplied .09*18.55=1.6695 then added 1.6695+18.55=20.2195 then rounded to 20.22

OpenStudy (tkhunny):

You can also do the addition first. (1 + 0.09)*18.55 = (1.09)*18.55 = 20.2195 Good work!

OpenStudy (anonymous):

thanks